OpenStudy (qwertyzxv):

3. What are the solutions of the system? y = x + 5 y = x^2 - x + 2 4. What are the solutions of the system? y = x^2 - 2x - 8 y = x + 2

OpenStudy (fortytherapper):

Since both are already in y =, set them equal to each other and solve for X. What do you get as your X for the first one?

OpenStudy (qwertyzxv):

i'm not sure, i'm pretty terrible at math

OpenStudy (fortytherapper):

Alright, lets set them equal X + 5 = X^2 - X + 2 When solving X, you wanna do the opposite operations until you get X by itself. So on the left, we have a positive 5. Since we do opposites, we would subtract 5 on both sides

OpenStudy (qwertyzxv):

i forgot to put up choices 3. What are the solutions of the system? y = x + 5 y = x^2 - x + 2 A. (3,14) and (-1,0) B. (3,8) and (-1,4) C. (3,8) and (-1,0) D. (-3,2) and (1,6) 4. What are the solutions of the system? y = x^2 - 2x - 8 y = x + 2 A. (5,7) and (-2,0) B. (5,3) and (-2,4) C. (-5,-3) and (2,4) D. (5,3) and (2-4)

OpenStudy (anonymous):

so both of these systems (number 3) are equal to y, meaning that they are equal to one another ex: 4+5=y 3+6=y 4+5=3+6 they both equal 9 in the same sense you equal both of these equations to one another y = x + 5 y = x^2 - x + 2 will be x + 5=x^2 - x + 2 from here solve (tell me what you get and i will confirm)

OpenStudy (anonymous):

and this process is used to answer number 4 as well

OpenStudy (qwertyzxv):

it has to be a solution like (0,0) i think i need a graph

OpenStudy (anonymous):

ok so after you set them equal to each other you will get the answer for x which you will simply have to plug back into the system and get y so your first answer will be the x and then plug in to get y (x,y)

OpenStudy (kikuo):

Is this Algebra 2?

OpenStudy (qwertyzxv):

no its algebra 1